OpenStudy (anonymous):

help medal and fan show work TIFFANY is filling her daughters pool with water from a hose.She can fill the pool at a rate of one tenth gallons/seconds. Create.Create at least three equivalent ratios that are associated with the rate.Use a double number line to show you work. for number one its a fraction one on top 10 bottom

OpenStudy (mathstudent55):

\(\dfrac{1}{10} = \dfrac{2}{20} = \dfrac{3}{30} \)

OpenStudy (anonymous):

not with math whith writting

OpenStudy (quickstudent):

I don't really understand what you need help with.

OpenStudy (anonymous):

not with math with writting

OpenStudy (mathstudent55):

One-tenth gallon per second is equivalent to two-tenths gallon in two seconds and is also equivalent to three-tenth gallons in three seconds.